Suppose you pay $0.30 to roll a fair 9-sided die with the understanding that you will get $0.70 back for rolling a 1, 2, or 3. Otherwise, you get no money back. What is your expected value of gain or loss?
Answer #1

Answers:

P(x=1,2 or3)=3/9=1/3

P(x=4,5,6,7,8 or 9)=5/9

Expected loss=(1/3)*0.70+(5/9)*0-0.30= -0.06666